‘n' identical cells are joined in series with two cells A and B with reverse polarities. Calculate the potential difference across the cell A and B .

Grade:12

1 Answers

Vikas TU
7 years ago
Asuming internal resistances be r and emf e for each cell joined in series.
then Total emf = ne
total resistance = nr
Total Potential Differnce,
(n-2)e – 2e = e(n-4)
V = ir + e
 = 2e(1-2/n)